OpenStudy (anonymous):

Anyone calculus here? Integrate 2pi(1/x)*sqrt(1 + 1/(x^4))?

OpenStudy (anonymous):

So to solve this question make the equation easier to solve (2(pi))/x + 2(pi)x^-4)/x = (2(pi))/x + 2(pi)x^-5 integrate function Use log rule for first fraction integral of 2(pi)/x = 2(pi)logx Then integrate normally with second fraction integral of 2(pi)x^-5= 2(pi)x^-4/-4 Add both integrals, therefore the final solution is; 2(pi)logx - (pi)/2x^4

OpenStudy (anonymous):

Also just in case you didn't notice or know about. I moved the x up to the numerator by make the power a negative e.g 1/x^4 = x^-4 Also x^-4/x = x^-5

OpenStudy (anonymous):

It's a long process..... i was tired....!! good question though

hartnn (hartnn):

you didn't consider sqrt ? \(\sqrt {1+1/x^4}\)

OpenStudy (anonymous):

yeah i checked that...

hartnn (hartnn):

i don't think that function can be integrated in terms of standard functions where did u get this question from ?

OpenStudy (anonymous):

it's an old IIT question.... !!

hartnn (hartnn):

hmm, i can think of a substitution \(\Large 1/x^2 = u \) need to work it out to check whether it works

OpenStudy (anonymous):

I thought same at first instance but it can be integrated...

OpenStudy (xapproachesinfinity):

Check by part, there's a bit of work required here I guess

OpenStudy (anonymous):

That's all i can get.. please verify if there's a mistake or need further calculation.

OpenStudy (xapproachesinfinity):

More you can try u substitution for the entire root thingy. It should work nice I guess

OpenStudy (anonymous):

calculation is lengthy enough..and after understanding it..it's hell lot of work.

OpenStudy (xapproachesinfinity):

Well you won't have that much of work compared to by parts

OpenStudy (anonymous):

ok i will try.. thanks!

OpenStudy (xapproachesinfinity):

Anytime